OpenStudy (anonymous):

The cost of labor varies jointly as the number of workers and the number of days they work. If it costs $3,136 to have 28 people work for 8 days, how much will it cost to have 20 people work for 16 days?

OpenStudy (anonymous):

See 28 persons in 8 days earn $3,136 Then, 1 person in 8 days would earn 3136/28= $112 Therefore 1person in 16 days would earn = 2( 1 person in 8 days) 1person in 16 days would earn = 2 (112) = $224 Therefore 20 persons in 16 days would earn = 20*(1person in 16 days earn) = 20(224) 20 persons in 16 days would earn = $4480

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

Welcome :)